OpenStudy (anonymous):

√(x) + √(x-7) = (21/√(x-7)) Solve the equation.

OpenStudy (anonymous):

\[\sqrt{x} + \sqrt{x-7} = \frac{ 21 }{ \sqrt{x-7} }\]

OpenStudy (tkhunny):

1) Write down \(x \ne 7\) 2) Multiply by \(\sqrt{x-7}\)